• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 10:48:00 on August 14, 2006

    Entry number 179836

    keyword=Beam Time Accounting
    Hall A

    08/13/06 0000-0800
    Experiment #: LEDEX
    Shift Leader: M. Olson
    Shift Workers: R. Michaels

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0.210.220.000.000.000.000.000.000.42
    Beam Available
    but Not in Use
    0.750.690.950.930.980.990.860.967.12
    Beam Not Available
    or Unacceptable
    0.050.090.040.070.020.010.140.040.46
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    0.420.250.000.000.000.000.000.000.67
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.580.751.001.001.001.001.001.007.33
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.580.751.001.001.001.001.001.007.33
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: Until 1:30 tried to use harps for calibration runs, but failure of Harp at H03B prevented that. Shift down for cooldown and refill of L-HRS cryogenic system.

    Data Entry by: R. Feuerbach